# Mixtures of Neural Network Experts with Application to Phytoplankton Flow Cytometry Data

### v1

@@ 6 Discussion
There is room for further model development. Our method assumes clusters are Gaussian, yet cell clusters in cytograms often exhibit skewness and heavy tails. A Bayesian semiparametric extension may provide a better fit to the data with the added benefits of uncertainty quantification and hyperparameter tuning without resampling procedures or cross-validation. Another area for future work relates to the hardware limitations of flow cytometers; the Gradients 2 dataset contains some observations which are censored at SeaFlow’s measurement limits. Proper handling of censored data may strengthen model validity and accuracy. Finally, it would be valuable to extend our application to a larger collection of cruise datasets exploring an even wider range of locations and times. Mapping the spatial distributions and average optical properties of phytoplankton regionally or globally based on in situ data, uncovering temporal trends, and identifying the environmental conditions which are most predictive of phytoplankton behavior, are ambitious yet exciting possibilities.
@@ Acknowledgements
This work was supported through grants by the Simons Collaboration on Computational Biogeochemical Modeling of Marine Ecosystems/CBIOMES (Grant ID: 1195553 to Sangwon Hyun) and the Simons Microbial Oceanography Project Award (Grant ID: 574495 to Francois Ribalet) and the National Science Foundation award NCSE-2215169. The authors also acknowledge the Hummingbird Computational Cluster at the University of California, Santa Cruz, for providing computing resources that have contributed to the research results reported within this publication. URL: https://hummingbird.ucsc.edu/.
@@
Figure 8: Principal component 2 and its most important constituent covariates over time: iron (Fe), sea level anomaly (SLA) and wind speed.
Figure 9: All nine principal components used in our data analysis.
@@
Figure 10: Estimated means and probabilities over time of Synechococcus for each model.
@@
Figure 11: Estimated means and probabilities over time of one of the PicoEukaryote clusters (Pico 1) for each model.
@@
Figure 12: Estimated means and probabilities over time of one of the PicoEukaryote clusters (Pico 2) for each model.
Supplement C: Individual Conditional Expectations
@@
Figure 13: Partial response curves of each response variable for each species (Pro, Syn, Pico1, and Pico2) as functions of the second principal component.
@@
Figure 14: Partial response curves of each response variable for each species as functions of the second principal component.

### v2

@@ 6 Discussion
There is room for further model development. Our method assumes clusters are Gaussian, yet cell clusters in cytograms often exhibit skewness and heavy tails. For this reason, it is worthwhile to explore modifications with skew Normal or skew $t$ kernels as well as nonparametric kernel specifications, such as infinite mixtures. In fact, a Bayesian semiparametric extension may provide a better fit to the data with the added benefits of uncertainty quantification and hyperparameter tuning without resampling procedures or cross-validation. Another area for future work relates to the hardware limitations of flow cytometers; the Gradients 2 dataset contains some observations which are censored at SeaFlow’s measurement limits. Proper handling of censored data may strengthen model validity and accuracy. Additionally, in the oceanographic setting, the responses are trivariate, whereas, in biomedical settings, the dimension is often much larger. This may introduce numerical instabilities and computational bottlenecks in estimating the covariance matrix of each Gaussian mixture component. In these cases, sparse precision matrix estimation techniques [31, 13] may be necessary. We also note that PCA does not induce sparsity in the original variables. The use of a method such as sparse PCA [26, 49] or group lasso [48] may provide sparser and therefore more easily interpretable latent factors. Our experiment in Supplement F also hints that some latent factors may contribute to little variability in the covariate space, but predict the response data well. This motivates an extension of our model which learns supervised latent factors, similar in spirit to partial least squares regression [46]. The correlated design and the potential for bias in the lasso estimator also warrants consideration of alternative coefficient penalties, including the trace lasso [15], adaptive lasso [50], and smoothly clipped absolute deviation penalty [10]. Finally, it would be valuable to extend our application to a larger collection of cruise datasets exploring an even wider range of locations and times. Mapping the spatial distributions and average optical properties of phytoplankton regionally or globally based on in situ data, uncovering temporal trends, and identifying the environmental conditions which are most predictive of phytoplankton behavior, are ambitious yet exciting possibilities.
@@ Acknowledgements
The authors acknowledge the Hummingbird Computational Cluster at the University of California, Santa Cruz, for providing computing resources that have contributed to the research results reported within this publication. URL: https://hummingbird.ucsc.edu/.
@@
Figure S1: Principal component 2 and its most important constituent covariates over time: iron (Fe), sea level anomaly (SLA) and wind speed.
Figure S2: All nine principal components used in our data analysis.
@@
Figure S3: Estimated means and probabilities over time of Synechococcus for each model.
@@
Figure S4: Estimated means and probabilities over time of one of the PicoEukaryote clusters (Pico 1) for each model.
@@
Figure S5: Estimated means and probabilities over time of one of the PicoEukaryote clusters (Pico 2) for each model.
Supplement C: Partial Response Curves
@@
Figure S6: Partial response curves of each response variable for each species (Pro, Syn, Pico1, and Pico2) as functions of the second principal component.
@@
Figure S7: Partial response curves of each response variable for each species as functions of the second principal component.
Supplement D: Simulation with a Multivariate Response
We conduct a simple three-dimensional simulation study to demonstrate the ability of the proposed model to recover a mixture of multivariate Gaussian components. For all time points $t$, we simulate $n_{t}=1,000$ particles and bin them using $D=50$ bins in each dimension. The two clusters are generated with the following parameters:
[…]
$\boldsymbol{R}_{1}$ $=\begin{pmatrix}1&0.5&0\\ 0.5&1&0\\ 0&0&1\end{pmatrix}$
$\boldsymbol{R}_{2}$ $=\begin{pmatrix}1&-0.7&0\\ -0.7&1&0.3\\ 0&0.3&1\end{pmatrix}$
[…]
Figure S8 overlays the estimated clusters on the data at two representative time points. Although the clusters exhibit substantial overlap at $t=67$, the proposed model accurately recovers their locations and covariance structure by borrowing information from time points with greater separation, such as $t=192$. Figure S9 shows that the cluster means and mixing probabilities are recovered with near-perfect accuracy.
Figure S8: Three-dimensional simulated cytograms at two time points, $t=67$ and $t=192$, with nonlinear model estimates overlaid in red. The data in the background is binned, with the intensity of blue hue proportional to the total number of particles in each bin. The solid points mark the cluster means, and the size of the points are proportional to the estimated cluster probabilities. The ellipses represent 2-dimensional projections of symmetric cluster-conditional 95% probability sets of each cluster.
Figure S9: Mean and probability estimates over time for two three-dimensional simulated clusters, compared with ground truth cluster parameters. The three dimensions of the cluster mean are displayed in the top-left, top-right, and bottom-left panels. The bottom-right panel compares the estimates of the mixing probability of the first cluster with the true probability. The estimated parameters for the first and second clusters are represented by orange and blue solid lines, respectively. True parameter values are represented by black dot-dashed lines.
Supplement E: Robustness to Hidden Layer Width and Randomness
To assess the robustness of our approach to the choice of hidden layer width (number of hidden nodes), we repeat the experiment shown in Figure 2 of the main text, for hidden widths $n_{h}=35,70,105,140$, and 175. As shown in Figure S10, the resulting estimates are nearly indistinguishable across these settings. However, the estimated mean of cluster 2 deviates from the true mean as the hidden width moves farther from $n_{h}=70$. This choice of $n_{h}=70$ for the main results in this article therefore represents a reasonable compromise between flexibility and stability in estimating nonlinear functions.
Resampling the hidden weights may introduce small variations in the estimates, but the method is generally robust to this randomness. In Figure S11 we display mean estimates averaged over five models with $n_{h}=70$ hidden widths, each fit using weights generated from different random number generator seeds. Overall, the estimates are very similar across models. Some variability is observed during the interval $t\in[10,150]$, where the amplitude of the oscillating mean changes. The model’s ability to capture nonlinear behavior therefore depends to some extent upon the particular draw of the hidden layer weights.
To increase the chances of obtaining favorable weights, two strategies may be considered. One could estimate multiple models with different random draws of the hidden weights and select the model with the lowest out-of-sample negative log-pseudolikelihood. Alternatively, one could estimate a single model with a large number of hidden nodes and allow the variable selection property of the lasso penalties to retain the hidden weights that contribute most to predictive performance. We prefer the latter approach, since it reduces the number of hyperparameters to tune and is less computationally demanding, requiring estimation of only a single model.
Figure S10: Estimated nonlinear models on simulated 1-dimensional data with a variety of hidden layer widths. All panels display in the background the simulated and binned data $y_{b}^{(t)}$ over time $t$, generated from the “Interaction Mean” model, with signal size $\Delta=0.40384$. The greyscale color of the bin $b$ at time $t$ is proportional to the biomass $c_{b}^{(t)}$. Overlaid are the estimated parameters $\hat{\boldsymbol{\mu}}_{k,t}$ over time $t$ from the nonlinear model as solid colored lines. These lines’ thickness at time $t$ is proportional to the estimated cluster probability $\hat{\pi}_{k,t}$. The thin colored lines are the bounds of symmetric 95% pointwise conditional probability regions of each Gaussian. The black dashed lines represent the true cluster means, and the black dotted lines represent the true 95% central probability regions.
Figure S11: Average of 5 estimated nonlinear models with randomly resampled hidden layer weights on simulated 1-dimensional data. All panels display in the background the simulated and binned data $y_{b}^{(t)}$ over time $t$, generated from the “Interaction Mean” model, with signal size $\Delta=0.40384$. The greyscale color of the bin $b$ at time $t$ is proportional to the biomass $c_{b}^{(t)}$. These lines’ thickness at time $t$ is constant with respect to $\hat{\pi}_{k,t}$ in order to avoid obscuring the uncertainty intervals. Approximate 95% confidence intervals around the model average—assuming a $t$-distribution with 4 degrees of freedom—are highlighted in purple.
Supplement F: Number of Principal Components
Here, we evaluate the stability of our model to the number of principal components used as regressors. We repeat the experiment in Section 5 of the main text, with $q=9,18,27$, and $37$ principal components and with the number of hidden nodes fixed at $n_{h}=70$. Since the principal components are used to improve interpretation, rather than model fit, we evaluate stability based on the variability of the partial response curves over different values of $q$, as displayed in Figure S12. There are some discrepancies in the estimated PicoEukaryote cluster 2 relative abundances at large values of PC1, which shows the potential usefulness of the trailing principal components beyond the first handful. But overall, the partial response curves change very little as a function of $q$. Additionally, the average 5-fold cross-validation score across the four models was 3.240 with a standard error of 0.016, indicating nearly identical predictive performance.
Figure S12: Partial responses of Prochlorococcus diameter, Synechococcus phycoerythrin fluorescence, and the second PicoEukaryote cluster’s relative abundance as functions of the first principal component. A separate curve is plotted for each number of principal components used as regressors in the nonlinear model. Predictions from the models using 9, 18, 27, and 37 principal components are represented by the solid, short-dashed, dot-dashed, and long-dashed lines, respectively.
Supplement G: Effect of Likelihood Biomass Weights
In order to avoid underestimating the prevalence of massive but rare particles, we raise the density of each particle $\boldsymbol{y}_{b}^{(t)}$ to a power proportional to the total biomass within that bin, $\boldsymbol{c}_{b}^{(t)}$, as shown in Equation (3) in the main text. Here, we compare this approach to the binned counts representation of the data, in which the power is simply the number of particles within the bin. The likelihood resulting from the binned counts representation is proportional to the likelihood of a discretized sample of the data with bin $b$ observed $c_{b}^{(t)}$ times. It is therefore a proper likelihood.
To showcase the effect of using biomass weights on parameter estimation, we cross-validate and refit our proposed model on the SeaFlow cruise data as in Section 5 of the main text twice: once with the binned biomass representation, and again with the binned counts representation.
Expert annotation by manual gating shows that there are four photosynthetic clusters: Prochlorococcus, Synechococcus, and PicoEukaryote 1 and 2. As we see in the left panel of Figure S13 using a pseudolikelihood (bins being upweighted by their biomass) recovers these expert-annotated clusters much more precisely, with each 95% high density region tightly enclosing high-biomass regions of cytogram space and reasonable probability estimates based on the amount of biomass seen within each cluster.
On the other hand, the right panel of Figure S13 clearly shows that omitting the biomass weights leads the model estimate to have many clusters whose cell diameter and chlorophyll are small. There are 6 estimated clusters tightly bunched around the general region where Prochlorococcus is located, with many of them highly overlapping. Only one potential large Gaussian cluster captures where PicoEukaryotes generally are. This cluster, as well as Synechococcus, have means and 95% high density regions which do not concentrate precisely around high biomass regions. This shows that the likelihood-based model which is based on the relative counts of the particles, is inferior in its ability to cluster populations compared to the pseudolikelihood-based model that is based on the relative biomass contribution of each subpopulation – the latter being more scientifically relevant.
Figure S13: Cytogram in two dimensions, diameter and chlorophyll, at time $t=33$ (2017-07-02 08:00-09:$0$), with model estimates overlaid in red (marking the important clusters) or grey (the rest). The data in the background is binned, with the intensity of blue hue proportional to the total biomass in each bin. The solid points mark the cluster means, and the size of the points are proportional to the estimated cluster probabilities. The ellipses represent 2-dimensional projections of symmetric cluster-conditional 95% probability sets of each cluster. The model estimates on the left are obtained using scaled particle biomasses as likelihood weights, while the estimates on the right are obtained use weights proportional to the number of particles within each bin.